Real Analysis

Math 310  MWF 10:30-11:25am
Fall 1996  127 Fowler

Ron Buckmire ∼ 402 Fowler ∼ 259-2536 ∼ ron@oxy.edu ∼ Mon 1:30-3:30pm Wed & Thu 1:30-2:30pm

(09/18/96)

SUMMARY Learning to write and read proofs

READING §8

\underline{Warm-Up} (5-10 minutes)

We will review your formal proofs of lim\frac{n + 1}{2n + 3} = 1/2

THINKING ABOUT PROOFS Today we are going to be doing more examples of formal proofs of the convergence of sequences. We shall be using this as a background for developing the skill of reading and writing formal mathematics. Example 2 (p. 29)

Prove that lim \frac{3n + 1}{7n − 4} = 3/7 Example 3 (p. 29)

Prove that lim \frac{4n^{3} + 3n}{n^{3} − 6} = 4

Example 4 (p. 29) Show that a_{n} = (−1)^{n} does not converge.

Weekly Homework Assignment (Due Friday 9/20/96)) §8, page 32: 8.1(a),(c), 8.2(a),(e) and 8.7(b)